# More Modality, More AI: Exploring Design Opportunities of AI-Based Multi-modal Remote Monitoring Technologies for Early Detection of Mental Health Sequelae in Youth Concussion Patients

### v1

@@ 8. Limitations and Future Work
This study identifies the challenges and needs of concussion clinicians when detecting mental health sequelae among youth concussion patients. It also presents design considerations and a proposed UI to support clinicians in this process. However, the study has several limitations. First, the sample size of our participants was limited. Six concussion clinicians participated in both the formative study and heuristic evaluation. However, recruiting these experts was challenging due to their demanding workloads (Wang et al., 2021; Jin et al., 2020) and the specialized expertise required to diagnose youth concussion patients with mental health sequelae. As a result, we conducted our study with highly relevant experts in this specific domain. Previous research has also employed a similar number of expert participants in related studies (Cai et al., 2019b; Beede et al., 2020; Jacobs et al., 2021; Yang et al., 2024; Zhang et al., 2024). Additionally, our recruited clinicians came from different hospitals and regions across the United States, enhancing sample diversity. Future research should expand the participant pool to include a broader range of clinicians across different geographic areas to improve the generalizability of the findings. Second, the final UI design was designed based on feedback from concussion clinicians during the heuristic evaluation. However, the final design has not yet been evaluated by concussion clinicians to assess its effectiveness and usability. Future research should incorporate participatory design methodologies to further refine the UI, ensuring it aligns more closely with clinicians’ workflows and needs. Third, our study focused on UI design rather than developing an interactive prototype. The primary goal of our UI design was to explore the feasibility and provide insights for further design refinements before actual development, thereby saving resources and time for both researchers and clinicians. Future research should investigate clinicians’ pain points and needs when interacting with a functional system developed. Finally, our study scenario specifically examined the detection of mental health sequelae after concussions among youth and explored the potential of AI-RPM in this context. Future studies could explore the broader application of AI-RPM in various clinical settings and understand how AI-driven systems can assist clinicians in early detection and intervention across different medical domains.
@@ 9. Conclusion
We explored the challenges and needs faced by concussion clinicians related to mental health issues in youth concussion patients and proposed design considerations for the use of AI-RPM technologies in this high-risk, high-stakes context. Our study makes two key contributions to the CSCW field. First, we identified three core challenges clinicians encounter when collecting and monitoring mental health-related information and making mental health-related decisions: (1) collecting patients’ mental health information outside the clinic; (2) communicating effectively with patients and their family members; and (3) assessing patients’ compliance with clinical recommendations. Second, we proposed a series of opportunities for AI-RPM technologies to address these challenges, as well as design considerations for AI-powered systems to collect multimodal, non-clinical data that supports clinicians in making informed mental health-related decisions. To our knowledge, this is the first study in the CSCW field to focus on the intersection of concussion clinicians and youth concussion patients in the context of mental health-related decision-making, providing a foundation for designing AI-RPM systems in similar scenarios.
@@ A Recent Experience
(1) Do you use any existing tools or features in EHR (e.g., EPIC) in the youth patients’ mental health decision-making? If so, what are they and how do you like them?
Appendix B APPENDIX: INTERVIEW SCRIPTS FOR DESIGN INTERFACE EVALUATION

### v2

@@ 8. Limitations and Future Work
Our work is not without limitations. First, the sample size of our participants was limited, with six concussion clinicians participating in both the formative and evaluation studies. Recruiting these experts was particularly challenging due to their demanding workloads (Wang et al., 2021; Jin et al., 2020) and the specialized expertise required to treat youth concussion patients with mental health issues As a result, we recruited six highly relevant experts in this specific domain. Despite the limited number, our inductive thematic analysis reached thematic saturation. Moreover, previous research employed a similar number of expert participants in related studies (Cai et al., 2019b; Beede et al., 2020; Jacobs et al., 2021; Yang et al., 2024; Zhang et al., 2024), which supports the appropriateness of our sample size. Future research should include a broader range of concussion clinicians across locations for greater generalizability of the findings.
Secondly, the refined AI-RPM system has not yet been further evaluated by concussion clinicians to assess its effectiveness and usability. In future research, we plan to incorporate participatory design methodologies (Muller and Kuhn, 1993) to revise the system design, ensuring it aligns more closely with clinicians’ workflows and needs. Moreover, our study focused on system design rather than developing an interactive prototype. The primary goal of our system design was to explore the feasibility and provide insights for further design refinements before actual development, thereby saving resources and time for both researchers and clinicians. Future research should investigate clinicians’ pain points and needs when interacting with a functional system developed.
Finally, we want to highlight the potential of applying AI-RPM systems in varied healthcare settings. Several studies have explored remote monitoring in the healthcare domain.  Wyche et al. (2024) utilized mHealth to track the health conditions of type 1 diabetes among youth.  Seals et al. (2022) leveraged wearables to detect gait impairment, monitor patients’ responses to treatment, and visualize patient data. We believe that AI-RPM technologies have the potential to be leveraged and extended to these scenarios. Future work should expand AI-RPM technologies to diverse healthcare settings, enabling timely interventions and improving care quality.
@@ 9. Conclusion
We conducted a formative study with six concussion clinicians to understand the challenges and needs during their clinical practice with youth concussion patients, particularly with respect to the patient mental health sequelae. Then, we derived a suite of design considerations with the use of AI-RPM technologies. Here are three design considerations. Firstly, we recommend the use of wearables to remotely monitor patients’ sleep and physical activity data, helping assess treatment compliance and inform mental health-related decisions. Secondly, we propose LLM-powered CAs to enhance mental health self-reporting, which facilitates private, natural interactions with patients and summarizes key information for clinicians. Thirdly, we suggest leveraging AI risk prediction in detecting concealed or worsening mental health sequelae. Finally we delivered a clinician-facing system design as our final artifact. To our knowledge, this is the first study in the CSCW field to focus on the intersection of health, RPM, and human-AI decision-making, which provides a foundation for designing AI-RPM systems in varies medical scenarios.
@@ A Recent Experience
(1) Do you use any existing tools or features in EHR (e.g., EPIC) in the youth patients’ mental health decision-making? If so, what are they, and how do you like them?
Appendix B APPENDIX: INTERVIEW SCRIPTS FOR EVALUATION STUDY